# Heterogeneous RBCs via deep multi-agent reinforcement learning

### v1

@@ 5. Conclusions
One obvious limitation of our framework is the computational cost required for accurately training the RL agents, with multi-agent training runs requiring up to a few hours to terminate successfully. However, easier access to vectorised computation on GPUs is likely to greatly mitigate this shortcoming in the near future. To move in this direction, it would be important to implement the MARL environment described here in a vectorised style in order to fully leverage GPU acceleration, for instance using specialised software Rutherford et al. (2024). The GPU acceleration (which can easily exceed a factor of 20) would allow much faster and more robust training, hence greatly facilitating the usability of MARL-BC. Furthermore, it would also allow for quicker model prototyping and extensions, and for the possibility of representing an even larger and more varied set of agents with RL.
While this study focused primarily on the description of MARL-BC, its limit cases, and its capabilities, future work could investigate the use of the framework to study specific economic problems. Natural issues could be related to economic inequality, but also to the economic and financial consequences of asymmetric changes in labour productivity, such as those potentially deriving from the spreading of AI tools in the workplace.
@@
By showing how a model of multiple interacting agents (akin to an ABM) can give rise to heterogeneous GE results, our work marks a step towards a synthesis of these often opposed methodologies. Furthermore, by showing how a MARL approach can recover and extend classic results from the economics literature, our work has a significant potential to foster communication and exchanges between communities in computer science and economics that are interested in RL modelling, but that have not previously found a common ground.
Acknowledgments
@@
We would like to thank Alessandro Moro, Valerio Astuti, Valerio Nispi Landi, Sergio Santoro (Banca d’Italia), Doyne Farmer (Oxford University), and Sara Casella (LUISS University) for useful feedback on this work.
The views and opinions expressed in this paper are those of the authors and do not necessarily reflect the official policy or position of Banca d’Italia.

### v2

@@ 5. Conclusions
One obvious limitation of our framework is the computational cost, with accurate training of RL agents requiring up to a few hours to terminate successfully. However, the use of vectorised GPU computation and of more modern MARL schemes is likely to greatly mitigate this shortcoming in the near future. Specifically, by implementing the MARL environment described in a GPU-compatible form, for instance using specialised software flair2024jaxmarl, would allow one to fully leverage GPU acceleration (which can easily exceed a factor of 20). Furthermore, while the independent learning scheme used here has the advantage of being simple to understand and implement, more sophisticated MARL approaches could improve learning stability and speed albrecht2024multi. This would allow much faster and more robust training, hence greatly facilitating usability, prototyping and extension of MARL-BC.
@@
While this study focused primarily on the description of MARL-BC, its limit cases, and its capabilities, future work could investigate the use of the framework to study specific economic problems. Natural issues could be related to economic inequality, but also to the economic and financial consequences of asymmetric changes in labour productivity, such as those potentially deriving from the spreading of AI tools in the workplace. By showing how a model of multiple interacting agents (akin to an ABM) can give rise to heterogeneous GE results, our work marks a step towards a synthesis of these often opposed methodologies. Furthermore, by showing how a MARL approach can recover and extend classic results from the economics literature, our work has significant potential to foster communication and exchanges between communities in computer science and economics that are interested in RL modelling, but that have not previously found common ground.
{acks}
@@
We would like to thank Alessandro Moro, Valerio Astuti, Valerio Nispi Landi, Sergio Santoro (Banca d’Italia), Doyne Farmer and José Morán (Oxford University), Marta Grześkiewicz (Cambridge University) and Sara Casella (LUISS) for useful feedback, and Michele Colombi (Scuola Normale Superiore) for double-checking some of the results.
Appendix A Appendix
We provide here additional details and experimental results complementing the information reported in the main text. First, in Sec. A.1, we document the hyperparameter settings used for each reinforcement learning (RL) algorithm to improve reproducibility. Then, in Sec. A.2, we report extended quantitative results on the learning process for the multi-agent experiments already described in the main text to further illustrate learning convergence and, in turn, the robustness and scalability of MARL-BC to large population sizes.
A.1. Hyperparameters of the RL algorithms
In the interest of reproducibility, we provide here further details on the hyperparameters of the RL algorithms used. All experiments were implemented using the default architectures provided in the Stable Baselines 3 library (v2.5.0). PPO employs a shared multilayer perceptron for policy and value estimation, with two 64-unit tanh layers, following the standard on-policy actor–critic architecture schulman2017proximal. SAC uses separate actor and critic networks, each with two 256-unit ReLU layers, optimising a stochastic policy under an entropy-regularised objective haarnoja2018softactorcriticoffpolicymaximum. TD3 and DDPG adopt two-hidden-layer networks (400 and 300 units, ReLU) for both actor and critics; TD3 additionally trains two critics independently and delays policy updates to mitigate value overestimation fujimoto2018td3; lillicrap2016ddpg. Table 2 summarizes the main learning hyperparameters adopted across algorithms in the experiments.
Code availability. We repeat here that the code used to generate the results is available at https://github.com/fedegabriele/MARL-BC.
PPO
Parameter Value
learning rate 3e-4
n steps 2048
batch size 64
n epochs 10
gae lambda 0.95
clip range 0.2
ent coef 0.0
vf coef 0.5
max grad norm 0.5
SAC
Parameter Value
learning rate 3e-4
buffer size 1e6
learning starts 100
batch size 256
tau 0.005
ent coef learned
target entropy learned
TD3
Parameter Value
learning rate 1e-3
buffer size 1e6
learning starts 100
batch size 256
tau 0.005
policy delay 2
target policy noise 0.2
target noise clip 0.5
DDPG
Parameter Value
learning rate 1e-3
buffer size 1e6
learning starts 100
batch size 256
tau 0.005
Table 2. Hyperparameters for PPO, SAC, TD3, and DDPG. Main hyperparameters for the four reinforcement learning algorithms used—Proximal Policy Optimization (PPO), Soft Actor-Critic (SAC), Twin Delayed Deep Deterministic Policy Gradient (TD3), and Deep Deterministic Policy Gradient (DDPG)—as implemented in Stable-Baselines3.
A.2. Extra results
We report here additional experimental results complementing those described in Sections 4.2 and 4.3 of the main text.
Extra results related to section ‘4.2 Mean-field Krusell–Smith limit’.
Figure 7 shows the learning curves of the four RL algorithms for increasing agent numbers (from $n=10$ to $n=500$) in the mean-field Krusell–Smith limit of MARL-BC. The figure shows that SAC consistently achieves high evaluation rewards across all population sizes. Specifically, for relatively small agent populations $n\leq 20$, SAC clearly outperforms any competing scheme while, as $n$ increases, PPO starts achieving similar rewards, although at a significantly slower pace. TD3 and DDPG underperform for all population sizes, particularly in the large-$n$ regime where they fail to learn stably. Importantly, the asymptotic reward level of SAC remains largely independent of $n$, indicating that the aggregate behaviour of the learned policy converges to the same mean-field equilibrium regardless of the population size. This result is corroborated by the left panel of Figure 8, which directly shows the maximum reward achieved as a function of population size. The right panel of the same figure confirms that SAC can be trained effectively within a reasonable time frame even on very modest hardware, as already demonstrated in Figure 6 of the main text.
Figure 9 shows the evolution of a number of quantities during training in the mean-field Krusell–Smith limit experiment. We note that as training progresses, the mean sum of discounted rewards ($\mathcal{R}^{i}$, bottom right panel) progressively grows and stabilises only slightly below $\mathcal{R}^{i}\approx 100$. In turn, this induces a state with consumption $c^{i}\approx 2.7$, capital $k^{i}\approx 25$, interest rate $r\approx 0.03$, and income $y^{i}\approx 3.3$, and with a Gini index on capital in the range 0.170.40.17-0.4.
Figure 7. Learning curves for mean-field Krusell–Smith limit experiment. Each panel shows the evaluation reward as a function of training steps for a specific number of agents ($n$ = 10, 20, 50, 100, 200, 500). Agents are ex-ante identical as in the mean-field (Krusell–Smith) model. The four RL algorithms are compared. Solid lines represent the median reward across 8 independent runs, and the shaded region represents the range between 25th and 75th percentiles. All panels share the same y-axis scale to facilitate direct comparison.
Figure 8. MARL-BC scaling to hundreds of agents for the mean-field Krusell–Smith limit. The left panel reports the best evaluation rewards achieved by MARL-BC agents during training runs of $10^{5}$ per-agent updates, meaning $n\cdot 10^{5}$ steps in total. The right panel reports the time needed for the corresponding runs, on a single CPU machine. Both quantities are shown as a function of the number of agents ($n$) in the model. Solid lines and shaded areas report mean $\pm$ one standard deviation across the 8 training runs. SAC is observed to obtain comparatively high rewards, in reasonable time limits, even for the largest models considered.
Figure 9. Training dynamics of the mean-field Krusell–Smith limit experiment with $n=20$ SAC agents. From left to right, top to bottom, the six panels show: individual consumption ($c^{i}$), individual capital ($k^{i}$), interest rate ($r$), individual income ($y^{i}$), Gini index over capital, and the individual sum of discounted returns ($\mathcal{R}^{i}$). The individual quantities are computed by first averaging over the agents, and the lines show the mean and percentiles over eight training runs.
Extra results related to section ‘4.3 Greater heterogeneity with MARL-BC’.
Figure 10 shows the learning curves of the four RL algorithms for increasing agent numbers (from $n=9$ to $n=529$) in the RBC model with heterogeneous agents with capital and labour productivities taken from an equispaced two-dimensional grid. For small populations $n\leq 20$, SAC, TD3, and DDPG rapidly achieve high evaluation rewards, while PPO converges more slowly and to lower values. As the number of agents grows, SAC remains the a very reliable learner, TD3 and DDPG show increasing instability and variance for large $n$. PPO, despite slower learning, improves performance for increasing $n$ exhibiting stable convergence for the largest population sizes considered. Importantly, the best attainable reward gradually converges toward a limiting value, suggesting that the heterogeneous system approaches a well-defined aggregate equilibrium for sufficiently large agent populations. This result is already illustrated in the top panel of Figure 6 of the main text, which directly reports the best rewards achieved as a function of population size for this model.
Figure 10. Learning curves for experiment on the RBC with grid-heterogeneous agents. Each panel shows the evaluation reward as a function of training steps for a specific number of agents ($n$ = 9, 25,49, 100, 225, 529). Agents have ex-ante different capital and labour productivities taken from an equispaced $\sqrt{n}\times\sqrt{n}$ grid ranging from 0.98 to 1.02. The four RL algorithms are compared. Solid lines represent the median reward across 8 independent runs, and the shaded region represents the range between 25th and 75th percentiles. All panels share the same y-axis scale to facilitate direct comparison.